1. Which solution has the lowest pH?

A. 0.10 M HClO4

B. 0.10 M HClO

C. 0.10 M KOH

D. 0.10 M CH3NH2

2. Which solution has the highest pH?

A. 0.05 M Sr(OH)2

B. 0.10 M Sr(OH)2

C. 0.10 M NaOH

D. 0.15 M NaOH

3. What is the pH of 0.03 M acid solution? The Ka of the acid is 510-3?

A. 1.5

B. 1.9

C. 2.0

D. 2.3 

4. What is the pH of 0.10 M Ba(OH)2 solution?

A. 0.70

B. 1.00

C. 13.00

D. 13.30 

5. What is the [OH- ] of a solution with pH = 4.33?

A. 2.1104 M

B. 2.110-10 M

C. 4.7109 M

D. 4.710-5 M

6. A 0.025 M solution of compound X has pH = 2.15. Which of the following statements is true?

A. Compound X is a strong acid.

B. Compound X is a strong base.

C. Compound X is a weak acid.

D. Compound X is a weak base.

7. A base has % ionization of 0.3%. What is the pH of 0.05 M solution of the base?

A. 1.3

B. 3.8

C. 10.2

D. 12.7


1
Expert's answer
2021-04-14T06:12:05-0400

1- "0.10MHCLO_4" has the lowest PH.

2- Ph of base ="pOH=-log(OH)"

then pH "=14-pOH"


3-"[H^+]= {K_aC}^\dfrac{1}{2}"

then pH ="-log(H^+)"


4- "[OH^-]={K_bC}^\dfrac{1}{2}"


then pOH="-log(OH)"

Now pH ="14-pOH"